Car Parts Center recently announced that it will pay annual dividends at the end of the next three years of $2.20, $2.25 and $2.35 per share, respectively. After that it will stop paying dividends. You expect to sell the shares in Year 5 at a price of $55.00 per share. At a required return of 15% percent, what should this stock sell for today?
Answer #1

D7 fax =SUM(D2:06) А Cash Flow 1 Year 21 32 2.25 PV@15% =1/1.15 =C2/1.15 =C3/1.15 =C4/1.15 =C5/1.15 Present Value =B2*C2 =B3*

D7 E 1 Year A B в Cash Flow 1 $ 2.20 2 $ 2.25 3 $ 2.35 4 $ - 5 $ 55.00 fx =SUM(D2:06) с PV@15% Present Value 0.8696 $ 1.91 0.